# The CLC-UKET Dataset: Benchmarking Case Outcome Prediction for the UK Employment Tribunal

### v1

@@ 1 Introduction
In recent years, there has been great interest in adapting natural language processing techniques in the legal domain. One notable application is the prediction of outcomes for legal disputes in various jurisdictions (Xiao et al., 2018; Poudyal et al., 2020; Hwang et al., 2022; Henderson et al., 2022; Niklaus et al., 2023). However, the AI-based prediction of UK court decisions is still under-explored.
@@
    We presented human performance on the UKET outcome prediction task on CLC-UKETpred with the aim of setting a human performance reference to calibrate prediction models.
@@
The CLC-UKET dataset and the empirical explorations aim to supplement the standard CLC dataset and facilitate future research on employment-related dispute resolution in the UK legal system. We will make the CLC-UKET dataset available with the final publication of this paper.
@@ 2 UK Legal System and UKET
The UK has a special category of judicial body, the employment tribunals, which deal exclusively with employment disputes. The UKET is one of the three largest tribunals in the greater tribunals system (Judicial Office, 2016). The UKET aims to provide a procedure which is easily accessible, informal, speedy and inexpensive (BEIS, 2020, p. 23). Claimants must comply with procedural and substantive requirements to be successful. For instance, claimants must submit their claims on time, comply with the orders of the tribunal, present required evidence or information in a timely manner, and avoid scandalous, unreasonable or vexatious conduct (which makes a fair trial impossible). These are usually considered as procedural requirements. Claimants must also comply with the substantial requirements of the rules supporting the claims. For example, in order to be successful with a discrimination claim on grounds of disability, the claimant must prove their status as an employee, demonstrate their disability and show that they faced discrimination, which are considered as substantial requirements of the case. These procedural and substantial requirements are not necessarily determined at one final hearing or included in one final judgment. Instead, they may be iteratively decided at different stages, which can result in multiple decisions.
@@
The employee (claimant) and the employer (respondent) submit their claims and responses, respectively, through a standardised form (Rules 8 and 16 of the Employment Tribunals Rules of Procedure 2013, hereinafter referred to as Rules ). The tribunal considers these forms and may dismiss a claim for procedural or substantial reasons, e.g., for lack of jurisdiction or for lack of any reasonable prospect of success (Rules, , r. 27). At any stage of the proceedings, the tribunal can determine a preliminary issue, make a procedural order (e.g., a deposit order or require the presentation of additional documents) or make a final decision (e.g., strike out the claim, Rules (, r. 37)). There may be multiple final hearings for different issues, for example, one hearing to determine whether or not a party is liable, another hearing to determine the remedy and another to determine who pays what costs (Rules, , r. 57). Each of these hearings results in a separate judgment, written out in a separate document. Finally, a party may request a reconsideration of a previous judgment, which will lead to another judgment (Rules, , r. 70). As a consequence, the resolution of a dispute may not be covered by one judgment only, but may be determined by iterative multiple decisions resulting in various case documents.
@@
Each decision is linked to one or multiple jurisdiction codes. In the case of UKET, there are 54 jurisdiction codes in total, which are used to identify the matter of disputes. By way of example, the jurisdiction code “unfair dismissal” is used when claimants argue that they have been unfairly dismissed. This jurisdiction code is often employed in addition to other jurisdiction codes, such as unlawful deduction from wages, redundancy, protective award, breach of contract, and working time regulations.
@@
In stark contrast with typical UK judgments, UKET decisions are relatively clearly structured, not only because there are no dissenting opinions, but also because there are specific rulings that set out which elements a judgment must contain (Rules, , r. 62(5)). Nevertheless, UKET judgments are not always consistent since there are no formal rules on the style to be used in drafting a decision. Most English judgments summarise their decisions in a paragraph, although this summary does not need to respect any particular form (d’État, 2012, p. 136). In the case of UKET, the summary is often found at the beginning of the judgment. However, judgments on multiple claims are sometimes divided into chapters, each analysing one claim containing the relevant decision. Also, while a judgment may contain an initial statement that the claimant is successful, it may not be clear which claim(s) this relates to in cases where there are multiple claims.
@@ 3.2 Legal Judgment Prediction
The advance of deep learning models alongside the development of large-scale legal datasets has greatly advanced the research on legal judgment prediction (LJP) (Xiao et al., 2018; Chalkidis et al., 2019; O’Sullivan and Beel, 2019; Ma et al., 2021; Chalkidis et al., 2023). A large number of datasets have been created for both civil law systems (Poudyal et al., 2020; Yamada et al., 2022) and common law systems (Caselaw Access Project, 2018; Henderson et al., 2022; Östling et al., 2023; Butler, 2024).
Facilitated by large-scale datasets, recent years have witnessed a surge in the application of deep learning models to LJP. Zhong et al. (2018) introduced TopJudge to address LJP using multi-task learning that combines three aspects: law articles, charges and terms of penalty. Another notable contribution is the work of Ma et al. (2021) where an end-to-end framework was built to predict dispute outcomes using multi-task supervision and multi-stage representation learning. To the best of our knowledge, the only notable LJP paper on UK law is Strickson and De La Iglesia (2020), which dates before the emergence of LLMs and is limited to the binary task of UK Supreme Court judges allowing or rejecting an appeal.
@@ 4.1 Case Preparation
The raw UKET subset of the CLC contains 52,339 cases in total, covering employment-related cases heard at the UKET from January 2011 to August 2023 (inclusive). [^1] The hearing dates of the cases in the UKET subset of the CLC range from 2011 to 2023, although some cases were submitted before 2011. After analysing these cases, we noticed that many cases only consist of one page as regards the tribunal decision. These cases are mostly the result of straightforward procedural decisions, for example when claimants withdraw their cases or respondents do not respond at all such that a default judgment is made. As these cases do not contain substantial information on the facts and the substantive reasons (de Faria et al., 2024), we excluded them at the case preparation step.
After this filtering step, we obtained a collection of 19,090 cases containing more than one page in their court decision files. For each case, we collected a list of metadata, including a unique case identifier used in the UKET records, date of filing, date of decision, place of hearing, judges, claimant(s), respondent(s) and appearances at the hearing. We also obtained jurisdiction codes for all cases from the UKET website. [^2] https://www.gov.uk/employment-tribunal-decisions. In legal contexts, a jurisdiction code typically refers to a numerical or alphanumeric code assigned to a specific legal jurisdiction, a certain subject matter or a geographic area. Legal jurisdictions are defined areas with a distinct set of laws and regulations. In the case of UKET, there are a total of 54 jurisdiction codes, which are used to identify the dispute matter. Each UKET case can be associated to multiple jurisdiction codes that indicate the categorical areas of the case. As an example, the code unfair dismissal is used when claimants argue that they have been unfairly dismissed (challenging the procedure of dismissal or the reason for dismissal) and submit a claim for payment of a certain sum, e.g., basic award, compensatory award (lack of notice pay and loss of earnings between a period) and injury to feelings award. A full list of jurisdiction codes in the UKET is presented in Appendix B.3.
@@ 4.2 LLM-aided Case Annotation
We applied an iterative development process to find the optimal prompt for the purpose of legal information extraction. The final prompt that we opted for yielded the best results in terms of the accuracy of information extracted, the adequacy of necessary information contained and the level of detail. The final prompt that we used for LLM-aided case annotation is presented in Appendix A.2. After automatic annotation, we obtained detailed annotations on important legal factors for 19,090 CLC-UKET cases, covering (1) facts, (2) claims, (3) references to legal statutes, (4) references to precedents, (5) general case outcomes, (6) general case outcomes labelled as “claimant wins”, “claimant loses”, “claimant partly wins” and “other”, (7) detailed orders and remedies and (8) reasons.
@@ 5.1 Task Definition
Given a set of facts and claims of a UKET case, the task of case outcome prediction aims to automatically generate an outcome label falling into one of the four categories of “claimant wins”, “claimant loses”, “claimant partly wins” and “other”. The facts and the claims are the judges’ summarisation of the statements provided by the claimant(s) and respondent(s) prior to or during a hearing.
@@
Note that there is a debate concerning the difference between the legal judgment prediction (LJP) task and the case outcome classification (COC) task (Medvedeva et al., 2021; Santosh et al., 2022; Medvedeva et al., 2023; Medvedeva and Mcbride, 2023). In this paper, we opted for the terminology “prediction” over “classification” as we deliberately excluded explicit information about case outcomes from the input of the prediction task, and only kept descriptions of facts and claims in the input. As such, this task focuses on predicting case outcomes based solely on information about case facts and applicants’ claims. Similarly, the legal experts predicting outcomes had only access to facts and claims.
@@ 5.2 Data Preparation for the Prediction Task
We tailored the CLC-UKETanno data to construct a case outcome prediction task for UKET. Three types of legal factors are needed for the prediction task, namely facts, claims and general case outcomes. The input to the prediction models is a sequence of fact statements concatenated with claim statements, in the form of “fact1, fact2, $\cdots$, factn [SEP] claim1, claim2, $\cdots$, claimm”. The target output of the prediction task is a general outcome label, which is a categorical variable labelling potential case outcomes as claimant wins, claimant partly wins, claimant loses and other.
@@ 6.1 Baseline Models
  • Zero-shot prediction. In this setting, the GPT-based models are directly asked to predict an outcome based on information about facts and claims of a case. No examples are provided for the models in the prompts.
@@ 6.2 Human Prediction
Two legal experts conducted the human prediction exercise. They are PhD candidates in Law with a focus on UK employment law. They were supervised by a professor of law based in the UK. Each test case in CLC-UKETpred was separately annotated by the two legal experts. We asked annotators to indicate what they think is the most likely case outcome after reading facts and claims of a case. They were also asked to indicate whether a prediction is of low confidence. Cases labelled with low confidence are usually cases that are hard to predict due to insufficient information contained in the given facts or claims or due to the intrinsic complexity of a case.
At the beginning of the annotation process, both annotators were provided with annotation guidelines (see Appendix D for details). The annotation guidelines are consistent with our overarching experimentation design for the prediction task. Annotators were asked to make their judgments separately, avoiding discussions amongst themselves. We emphasised that human predictions should be made based on the same facts and claims that prediction models were evaluated on. Annotators were required to not search for the cases they were annotating on the internet. [^4] However, annotators were free to research other information that might be helpful for the annotations, for example, information on the applicable law. Whenever questions regarding the implementation of the annotation arose during the annotation process, the annotators were provided with clarification by the supervisor.
After annotating, we obtained two independent sets of predicted case outcome labels for the 1,371 test cases. The Cohen’s Kappa score for all annotations is 0.421.
@@ 6.3 Results
Random 0.380 0.487 0.380 0.420
@@
Table 4: Evaluation results obtained by baseline models and human predictions when the labels “wins” and “partly wins” are aggregated. Random refers to random guess. Human refers to the averaged scores of the outcome labels predicted by two human experts. For GPT-3.5 and GPT-4, the variants that achieved the highest F-scores across relevant settings are presented.
@@ 7.1 Relevance of Scores
We would like to emphasise that the evaluation scores reported for the CLC-UKETpred prediction task are baseline results. Both the transformer-based and the LLM-based models could be improved further for the task at hand. For example, the latter could be further enhanced by incorporating retrieval-augmented generation (Lewis et al., 2020; Gao et al., 2023) or chain-of-thought (Wei et al., 2022; Diao et al., 2023; Kim et al., 2023). Similarly, human experts could invest more time and research into their predictions. Those interested in the legal domain are, therefore, encouraged to apply caution when drawing conclusions for legal practice.
Against this background, it is worth discussing a few patterns in the scores. First, both models and legal experts achieve higher recall than precision scores for “wins” and higher precision scores than recall scores for “loses”. Precision is a useful measure when the costs of a wrongly predicted positive are high. In a litigation context, this is the case when the costs of initiating litigation (e.g., fees for legal and other advisers, court fees, time and nerves invested etc) are high. Likewise, recall is a useful measure if the costs of missing a true positive are high. In the context of court proceedings, this is the case when the opportunity cost of not initiating likely successful litigation is high, for example, if the expected remedy has a high monetary value or otherwise has a high relevance for the potential claimant (e.g., for emotional reasons). Hence, it depends on the specific situation of a potential claimant whether precision or recall provides better guidance. Since the UKET currently does not charge fees and claimants can represent themselves (thereby saving costs), recall may be the preferable score if the claim matters to the potential claimant. Second, it is worth noting that the F-score of GPT-4juris2 for “partly wins” outperforms the human predictors. This may indicate the LLM’s ability to navigate more complex litigation, which involves multiple claims or multiple parties on either side.

### v2

@@ 1 Introduction
In recent years, there has been great interest in adopting natural language processing techniques in the legal domain. One notable application is the prediction of outcomes for legal disputes in various jurisdictions (Xiao et al., 2018; Poudyal et al., 2020; Hwang et al., 2022; Henderson et al., 2022; Niklaus et al., 2023). However, the AI-based prediction of UK court decisions is still under-explored.
@@
    We assessed human performance on the UKET outcome prediction task on CLC-UKETpred with the aim of setting a human performance reference to calibrate prediction models.
@@
The CLC-UKET dataset and the empirical explorations aim to supplement the standard CLC dataset and facilitate future research on employment-related dispute resolution in the UK legal system. We will make the CLC-UKET dataset available via the official CLC website [^1] The CLC website: https://www.cst.cam.ac.uk/research/srg/
projects/law.
@@ 2 UK Legal System and UKET
The UK has a special category of judicial body, the employment tribunals, which deal exclusively with employment disputes. The UKET is one of the three largest tribunals in the greater tribunals system (Judicial Office, 2016). The UKET aims to provide a procedure which is easily accessible, informal, speedy and inexpensive (BEIS, 2020, p. 23). The form of employment tribunal proceedings is adversarial rather than investigatory, as each party has to present and prove its case (Deakin et al., 2021). Claimants must comply with procedural and substantive requirements to be successful. For instance, claimants must submit their claims on time, comply with the orders of the tribunal, present required evidence or information in a timely manner, and avoid scandalous, unreasonable or vexatious conduct (which would make a fair trial impossible). These are usually considered as procedural requirements. Claimants must also comply with the substantive requirements of the rules supporting the claims. For example, in order to be successful with a discrimination claim on grounds of disability, the claimant must prove their status as an employee, demonstrate their disability and show that they faced discrimination, which are considered as substantive requirements of the case. These procedural and substantive requirements are not necessarily determined at one final hearing or included in one final judgment. Instead, they may be iteratively decided at different stages, which can result in multiple decisions.
@@
The employee (claimant) and the employer (respondent) submit their claims and responses, respectively, through a standardised form (Rules 8 and 16 of the Employment Tribunals Rules of Procedure 2013, hereinafter referred to as Rules ). The tribunal considers these forms and may dismiss a claim for procedural or substantive reasons, e.g., for lack of jurisdiction or for lack of any reasonable prospect of success (Rules, , r. 27). At any stage of the proceedings, the tribunal can determine a preliminary issue, make a procedural order (e.g., a deposit order or require the presentation of additional documents) or make a final decision (e.g., strike out the claim, Rules (, r. 37)). There may be multiple final hearings for different issues, for example, one hearing to determine whether a party is liable, another hearing to determine the remedy and another to determine the costs (Rules, , r. 57). Each of these hearings results in a separate judgment, written out in a separate document. Finally, a party may request a reconsideration of a previous judgment, which will lead to another judgment (Rules, , r. 70). As a consequence, the resolution of a dispute may not be covered by one judgment only, but may be determined by iterative multiple decisions resulting in various case documents.
@@
Each decision is linked to one or multiple jurisdiction codes. In the case of the UKET, there are 54 jurisdiction codes in total, which are used to identify the matter of disputes. By way of example, the jurisdiction code “unfair dismissal” is used when claimants argue that they have been unfairly dismissed. This jurisdiction code is often employed in addition to other jurisdiction codes, such as unlawful deduction from wages, redundancy, protective award, breach of contract and working time regulations.
@@
In stark contrast with typical UK judgments, UKET decisions are relatively clearly structured, not only because there are no dissenting opinions, but also because there are specific rulings that set out which elements a judgment must contain (Rules, , r. 62(5)). Nevertheless, UKET judgments are not always consistent since there are no formal rules on the style to be used in drafting a decision. Most English judgments summarise their decisions in a paragraph, although this summary does not need to respect any particular form (Conseil d’État, 2012, p. 136). In the case of the UKET, the summary is often found at the beginning of the judgment. However, judgments on multiple claims are sometimes divided into chapters, each analysing one claim containing the relevant decision. Also, whilst a judgment may contain an initial statement that the claimant is successful, it may not be clear which claim(s) this relates to in cases where there are multiple claims.
@@ 3.2 Legal Judgment Prediction
The advance of deep learning models alongside the development of large-scale legal datasets has greatly advanced the research on legal judgment prediction (LJP) (Xiao et al., 2018; Chalkidis et al., 2019; O’Sullivan and Beel, 2019; Ma et al., 2021; Chalkidis et al., 2023; Colombo et al., 2024). A large number of datasets have been created for both civil law systems (Poudyal et al., 2020; Yamada et al., 2022) and common law systems (Caselaw Access Project, 2018; Henderson et al., 2022; Östling et al., 2023; Butler, 2024).
Facilitated by large-scale datasets, there has been a surge in the application of deep learning models to LJP in recent years. Zhong et al. (2018) introduced TopJudge to address LJP using multi-task learning that combines three aspects: law articles, charges and terms of penalty. Another notable contribution is the work of Ma et al. (2021) where an end-to-end framework was built to predict dispute outcomes using multi-task supervision and multi-stage representation learning. To the best of our knowledge, the only notable LJP paper on UK law is Strickson and De La Iglesia (2020), which dates before the emergence of LLMs and is limited to the binary task of UK Supreme Court judges allowing or dismissing an appeal.
@@ 4.1 Case Preparation
The raw UKET subset of the CLC contains 52,339 cases in total, covering employment-related cases heard at the UKET from January 2011 to August 2023 (inclusive). [^2] The hearing dates of the cases in the UKET subset of the CLC range from 2011 to 2023, although some cases were submitted before 2011. After analysing these cases, we noticed that many cases only consist of one page as regards the tribunal decision. Based on the observations from de Faria et al. (2024), many of these cases involve straightforward procedural decisions, for example when claimants withdraw their cases or respondents do not respond at all such that a default judgment is made. As such cases do not contain substantial information on facts and substantive reasons, we excluded them at the case preparation step.
After this filtering step, we obtained a collection of 19,090 cases containing more than one page in their court decision files. For each case, we collected a list of metadata, including a unique case identifier used in the UKET records, date of filing, date of decision, place of hearing, judges, claimant(s), respondent(s) and appearances at the hearing. We also obtained jurisdiction codes for all cases from the UKET website. [^3] https://www.gov.uk/employment-tribunal-decisions. In legal contexts, a jurisdiction code typically refers to a numerical or alphanumeric code assigned to a specific legal jurisdiction, a certain subject matter or a geographic area. Legal jurisdictions are defined areas with a distinct set of laws and regulations. In the case of the UKET, there are a total of 54 jurisdiction codes, which are used to identify the dispute matter. Each UKET case can be associated to multiple jurisdiction codes that indicate the categorical areas of the case. As an example, the code unfair dismissal is used when claimants argue that they have been unfairly dismissed and submit a claim for payment of a certain sum, e.g., basic award, compensatory award (lack of notice pay and loss of earnings between a period) and injury to feelings award. A full list of jurisdiction codes in the UKET is presented in Appendix B.3.
@@ 4.2 LLM-aided Case Annotation
We applied an iterative development process to find the optimal prompt for the purpose of legal information extraction. The final prompt that we opted for yielded the best results in terms of the accuracy of information extracted, the adequacy of necessary information contained therein and the level of detail. The final prompt that we used for LLM-aided case annotation is presented in Appendix A.2. After automatic annotation, we obtained detailed annotations on important legal factors for 19,090 CLC-UKET cases, covering (1) facts, (2) claims, (3) references to legal statutes, (4) references to precedents, (5) general case outcomes, (6) general case outcomes labelled as “claimant wins”, “claimant loses”, “claimant partly wins” and “other”, (7) detailed orders and remedies and (8) reasons.
@@ 5.1 Task Definition
Given a set of facts and claims of a UKET case, the task of case outcome prediction aims to automatically generate an outcome label falling into one of the following four categories: “claimant wins”, “claimant loses”, “claimant partly wins” and “other”. The facts and the claims are the judges’ summarisation of the statements provided by the claimant(s) and respondent(s) prior to or during a hearing.
@@
Note that there is a debate concerning the difference between the legal judgment prediction (LJP) task and the case outcome classification (COC) task (Medvedeva et al., 2021; Santosh et al., 2022; Medvedeva et al., 2023; Medvedeva and Mcbride, 2023). In this paper, we opted for the terminology “prediction” over “classification” as we deliberately excluded explicit information about case outcomes from the input of the prediction task, and only kept descriptions of facts and claims in the input. As such, this task focuses on predicting case outcomes based solely on information about facts and claims. Similarly, the legal experts predicting outcomes had only access to facts and claims.
@@ 5.2 Data Preparation for the Prediction Task
We tailored the CLC-UKETanno data to construct a case outcome prediction task for the UKET. Three types of legal factors are needed for the prediction task, namely facts, claims and general case outcomes. The input to the prediction models is a sequence of fact statements concatenated with claim statements, in the form of “fact1, fact2, $\cdots$, factn [SEP] claim1, claim2, $\cdots$, claimm”. The target output of the prediction task is a general outcome label, which is a categorical variable labelling potential case outcomes as claimant wins, claimant partly wins, claimant loses and other.
@@ 6.1 Baseline Models
  • Zero-shot prediction. In this setting, the GPT-based models are directly asked to predict an outcome based on information about facts and claims of a case. No examples are provided to the models in the prompts.
@@ 6.2 Human Prediction
Two legal experts conducted the human prediction exercise. They are PhD candidates in Law with a focus on UK employment law. They were supervised by a professor of law based in the UK. Each test case in CLC-UKETpred was separately annotated by the two legal experts. We asked annotators to indicate what they think is the most likely case outcome after reading facts and claims of a case. They were also asked to indicate whether a prediction is of low confidence. Cases labelled with low confidence are usually cases that are hard to predict due to insufficient information contained in the given facts or claims or due to the intrinsic complexity of a case (in particular the claims raised).
At the beginning of the annotation process, both annotators were provided with annotation guidelines (see Appendix D for details). The annotation guidelines are consistent with our overarching experimentation design for the prediction task. Annotators were asked to make their judgments separately, avoiding discussions amongst themselves. We emphasised that human predictions should be made based on the same facts and claims that prediction models were evaluated on. Annotators were required not to search for the cases they were annotating on the internet. [^5] However, annotators were free to research other information that might be helpful for the annotations, for example, information on the applicable law. Whenever questions regarding the implementation of the annotation arose during the annotation process, the annotators were provided with clarification by the supervisor.
After annotating, we obtained two independent sets of predicted case outcome labels for the 1,371 test cases. The Cohen’s Kappa score for all annotations is 0.421666The Cohen’s Kappa score between two specialised legal experts ranges from 0.41 to 0.60 indicating moderate agreement, highlighting the inherent complexity in the UKET prediction task..
@@ 6.3 Results
BERT 0.443 0.619 0.443 0.421
T5 0.535 0.552 0.535 0.480
GPT-3.5juris5 0.455 0.488 0.455 0.451
GPT-4juris2 0.465 0.527 0.465 0.448
Human 0.477 0.507 0.477 0.448
Table 4: Evaluation results obtained by baseline models and human predictions on test cases which are considered as hard to predict by human experts.
Performance on low confidence cases. In the human prediction process described in Section 6.2, expert annotators were asked to explicitly indicate whether a case was difficult for them to predict based on the given facts and claims (i.e., a “low confidence” prediction). Using these annotations, we further analysed different baselines for cases that were considered difficult by the human experts. Comparing Table 4 with Table 2, it can be observed that human performance on predicting for the low confidence cases is significantly worse than for all cases, suggesting that human assessments of the difficulty level of the prediction task align well with the empirical results. Furthermore, all baseline models exhibit relatively lower scores when evaluated on the low confidence cases. This pattern indicates that cases that are more challenging for human experts are also more difficult for the models.
Baseline Accuracy Precision Recall F-score
@@
Table 5: Evaluation results obtained by baseline models and human predictions when the labels “wins” and “partly wins” are aggregated. Human refers to the averaged scores of the outcome labels predicted by two human experts. For GPT-3.5 and GPT-4, the variants that achieved the highest F-scores across relevant settings are presented.
@@ 7.1 Relevance of Scores
We would like to emphasise that the evaluation scores reported for the CLC-UKETpred prediction task are baseline results. Both the transformer-based and the LLM-based models could be improved further for the task at hand. For example, the latter could be further enhanced by incorporating retrieval-augmented generation (Lewis et al., 2020; Gao et al., 2023) or chain-of-thought (Wei et al., 2022; Diao et al., 2023; Kim et al., 2023). Similarly, human experts might achieve better predictions by investing more time and conducting further research. Those interested in the legal domain are, therefore, encouraged to apply caution when drawing conclusions for legal practice.
The prediction task has been designed from the perspective of the claimant. This perspective informs the outcomes “claimant wins”, “claimant loses”, “claimant partly wins” and “other”. This approach makes sense as it is first for the claimant to decide whether they apply for a decision of the Tribunal. Once the claimant has taken this first step, it is for the defendant to decide how they react to the claim. Whilst the outcome prediction for the claimant is also of relevance for the defendant, it should be noted that both models and human predictors achieve different scores depending on whether “wins” or “loses” is predicted.
Against this background, it is worth discussing a few patterns in the scores. First, both models and legal experts achieve higher recall than precision scores for “wins” and higher precision scores than recall scores for “loses”. Precision is a useful measure when the costs of a wrongly predicted positive are high. In a litigation context, this is the case when the costs of initiating litigation (e.g., fees for legal and other advisers, court fees, time and stress involved) are high. Likewise, recall is a useful measure if the costs of missing a true positive are high. In the context of court proceedings, this is the case when the opportunity cost of not initiating likely successful litigation is high, for example, if the expected remedy has a high monetary value or otherwise has a high relevance for the potential claimant (e.g., for emotional reasons). Hence, it depends on the specific situation of a potential claimant whether precision or recall provides better guidance. Since the UKET currently does not charge fees and claimants can represent themselves (thereby saving costs), recall may be the preferable score if the claim matters to the potential claimant. Second, it is worth noting that the F-score of GPT-4juris2 for “partly wins” outperforms the human predictors. This may indicate the LLM’s ability to navigate more complex litigation, which involves multiple claims or multiple parties on either side.